This discussion thread test the Learning Objective of 2-8, Compute the present value of a single amount and 6-7 Compute the present value of an ordinary annuity, and annuity due, and a deferred and annuity. Sally Hamilton has performed well as the chief financial officer of the Maxtech Computer Computer Company and has earned a bonus. She has a choice among the following three bonus plans: 1. A $50,00 cash bonus paid now. 2. a $10,000 annual cash bonus to be paid each year over the next six years, with the first $10,000 paid now. 3. A three-year$22,000 annual cash bonus with the first payment due three years from now. Required: Evaluate the three alternative bonus plans.

Sally can earn 6% annual return on her investment. Which would you select and why?
Answer #1

The option with the highest present value today should be selected

Option 1 = $50,000

Option 2 = 10,000 + 10,000*PVAF(6%, 5 years)

= 10,000 + 10,000*4.2124

= $52,124

Option 3 = Value after 2 years = 22000*PVAF(6%, 3 years)

= 22000*2.6730

= $58,806

Value today = 58,806/(1.06)^2

= $52,337.13

Hence, Option 3 should be chosen